OpenStudy (anonymous):

3. In your industrial oven, you bake two baking sheets with 10 scones each, two baking sheets with 24 cookies each, and one baking sheet with 2 scones and 16 cookies. a. Write an expression that illustrates the total cost of all baked goods in the scenario above using the variable s to represent the cost of scones, and the variable c to represent the cost of cookies. Simplify your expression by combining like terms. C=cost C=2(10s)+2(24c)+2s+16c (multiply like amounts in parenthesis) C= 20s + 48c + 2s + 16c (add like amounts c + s) C=22s +62c (simplify my expression) b. Imagi

OpenStudy (anonymous):

b. Imagine you have decided to price the scones at $2.35 each and the cookies at $0.99 each. How much total revenue would result from selling all the scones and cookies baked in the oven at one time? R=26(2.35)+62(.99) R=61.10+61.38 R=122.48 c. Yesterday your store earned $707.85 just from the sale of cookies. Write and solve an equation that represents how many cookies were sold. Cookies sold = 707.85/.99=715

OpenStudy (lilg132):

sorry should be 22 scones and 64 cookies how did you get 26 scones and 62 cookies?

OpenStudy (anonymous):

oops i added wrong, thanks for catching that...

OpenStudy (anonymous):

just corrected my answers..

OpenStudy (lilg132):

thats fine :)

OpenStudy (anonymous):

is c. correct?

OpenStudy (anonymous):

is the equation correct for c?

OpenStudy (lilg132):

CS = R/P CS = Cookies sold R = Revenue P = price so you should have 707.85/.99 = 715 CS

OpenStudy (anonymous):

awesome..

OpenStudy (anonymous):

and last one..will repost..and another medal, yay

OpenStudy (lilg132):

lol yes feel free